OpenStudy (anonymous):

Part A: Explain why the x coordinates of the points where the graphs of the equations y=4^x and y=2^x+2 interest are the solutions of the equation 4^x=2^x+2

OpenStudy (nikato):

Transitive If A=B and B=C, then A=C

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

Let's say we have two equations. \[y = 2 x\]and\[y = -2x\] If we want to find where their intersect on a graph, we need to understand how the graph relates to the equation. Remember how we graph simple equations. We let x = 1 then find y and then plot the point. Repeat for a few more values. The graph represents the value of y for each value of x of a given equation. If two equations intersect, that means that both equations have the same y-value given the same x-value. That means that we can set the two equations equal to each other, because they have the same y-value. From my earlier two equations, \[2x = - 2x\] Now, we can solve for the x-value that satisfies this condition of both equations having the same y-value. In this case, it is zero. Both graphs intersect at zero.

OpenStudy (anonymous):

Both graphs intersect at (0,0) to be more precise.

OpenStudy (anonymous):

@nikato is he correct?

OpenStudy (nikato):

Yes. Same process for solving your problem, but he was using his own example

OpenStudy (anonymous):

how did he find x?

OpenStudy (anonymous):

how did he find x though?

OpenStudy (anonymous):

My example was a bad one. You divide both sides by x and then are left with 2 = -2. This is arbitrary. If we subtract -2x from both sides, we get 2x + 2x = 0. The only value of x that satisfies the equation is zero. 4x = 0 --> x = 0 because 0/4 = 0. A better example might be \[y= x^2\]and\[y = 2x+3\] We set them equal to each other. \[x^3 = 2x+3\] Set the equation equal to zero\[x^2 - 2x- 3 = 0\] Now it looks like your run-of-the mill quadratic equation.

OpenStudy (anonymous):

You'll get two solutions because. 

OpenStudy (anonymous):

Can you explain with my equation?

OpenStudy (anonymous):

hang on let me try one moment please.

OpenStudy (anonymous):

Sure. \[y = 4^x\]and \[y = 2^x + 2\]

OpenStudy (anonymous):

Without spoiling your work, here is a neat visualization. http://www.wolframalpha.com/input/?i=y+%3D+4%5Ex+and+y+%3D+2%5Ex+%2B+2

OpenStudy (anonymous):

To find out the x coordinate you will do 4^x=2^x+2 and solve for x by subtracting 2^x from both sides leaving you with 2^x=2. Then, by subtracting 2 from both sides leaving you with 2^x-2=0. what do I do next?

OpenStudy (anonymous):

You cannot subtract 4^x and 2^x because they are dissimilar bases.

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

Getting these to the same base is easy in this case. Remember that\[2^2 = 4\] Therefore, we can rewrite the equation as\[4^x = (2^2)^x\]Recall that\[(x^y)^z = x^{yz}\] Therefore, \[2^{2x} = 2^x + 2\] Now they're in the same base.

OpenStudy (anonymous):

Solve as you previously attempted.

OpenStudy (anonymous):

so 2^2x-2^x-2=0?

OpenStudy (anonymous):

Yep!